OpenStudy (anonymous):

Hello Quick question on exponents. When asked to evaluate a.) (-3)^2 and b.) -3^2 I'm inclined to answer 9 for both. But, the answers given in the book they're taken from suggest a=9 and b=-9. Can someone explain why? Thanks

OpenStudy (anonymous):

the book is correct (-3)^2 = -3 * -3 = 9 - (3)^3 = - 3* 3 = -9 when you have a 'mixture; of add/subtract and multiply/ divide the multiply/divide is performed first

OpenStudy (anonymous):

well look at it this way: (-3)^2=(-3)(-3) a negative times a negative is a positive (thump thump) so (-3)(-3)=9 -3^2=-(3)(3) a positive times a positive is a positive so (3)(3)=(9) and -(3)(3)=-(9)

OpenStudy (anonymous):

the acronym BODMAS is a memory aid for the order of operations = brackets over divide/multiply over add/subtract

OpenStudy (anonymous):

I learned it Please Excuse My Dear Aunt Sally You start with Parentheses, then exponents, Multiplication, Division, Addition, and Subtraction

OpenStudy (amistre64):

(.)^2 is actually all one operator in this case; the exponent is a grouping symbol ...

OpenStudy (amistre64):

(..)^2 means, square whatever you have inside

OpenStudy (amistre64):

(-3)^2 means, -3 * -3 = 9 -(3)^2 mean, -(3*3) = -(9) = -9

OpenStudy (anonymous):

Thanks for all the replies lads What I'm unclear about is why -3^2 is not =9. I assumed squaring meant -3*-3 or multiplying the value by itself. I'm starting maths from scratch again trying to find out areas I've missed which have held me back. Order of operations was one such area but I must be missing something in exponent rules

OpenStudy (amistre64):

-3^2 has 2 symbolds to it; and the ^2 has inherent properties that the number it is directly attached to is its main focus

OpenStudy (amistre64):

-3^2 is the same as -(3)^2

OpenStudy (amistre64):

an exponent is a grouping symbol, is is just as groupy as brackets

OpenStudy (amistre64):

-3^2 reads; the opposite of the square of three.

OpenStudy (anonymous):

ok so it doesn't imply there's a calculation to be made, its just a representation?

OpenStudy (amistre64):

im not sure i understand your response

OpenStudy (amistre64):

it takes practice to get a good eye to read thru the math; its like learning a new language

OpenStudy (anonymous):

Well I suppose this makes the most sense to me -3^2 is the same as -(3)^2. If I stick to that as a rule and not ask why ;)

OpenStudy (amistre64):

the () are imlplied, thats all

OpenStudy (amistre64):

it like having an invisible 1 in front of a variable ... its just implied

OpenStudy (anonymous):

Well I suppose the fact that it's implied is what I missed somewhere down the road ;) Thanks to you and everyone else for yer time. Appreciate it

OpenStudy (anonymous):

i notice that some people have difficulty with simplifying something like -(x - 3) - they dont get how - 3 becomes + 3 after removing the brackets It helps if you put the invisible 1 into it : that is -1(x - 3) = -1*x - 1*-3 = -x + 3

OpenStudy (anonymous):

Got it, cheers mon